40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Feb 20 16:20:33 2021, gautam, Update, General, Housekeeping + PRMI char IPPOSdeat.pdfTTdrift.pdfPOXnPOY.pdfPRMI.pdf
    Reply  Sun Feb 21 20:51:21 2021, rana, Update, General, Housekeeping + PRMI char 
       Reply  Tue Feb 23 00:10:05 2021, gautam, Update, General, Demod char part 1 noPreamp.pdfwithPreamp.pdf
          Reply  Wed Feb 24 11:53:24 2021, gautam, Update, General, Demod char part 2 noises_noPreamp.pdfnoises_withPreamp.pdf
             Reply  Wed Feb 24 12:11:08 2021, gautam, Update, General, Demod char part 3 noPreamp.pdfwithPreamp.pdfnoPreamp_whitened.pdfwithPreamp_whitened.pdf
                Reply  Thu Feb 25 16:50:53 2021, gautam, Update, General, PRMI sensing matrix PRMI1f_noArmssensMat.pdf
                   Reply  Thu Feb 25 20:37:49 2021, gautam, Update, General, Setting modulation frequency and checking IMC offset modFreq.pdfIMC_offset.pdf
Message ID: 15845     Entry time: Thu Feb 25 20:37:49 2021     In reply to: 15844
Author: gautam 
Type: Update 
Category: General 
Subject: Setting modulation frequency and checking IMC offset 

The Marconi frequency was tuned by looking at 

  1. The ~3.68 MHz (= 3*f1 - fIMC) peak at the IMC servo error point, TP1A, and
  2. The ~25.8 MHz (= 5*f1 - fIMC) peak at the MC REFL PD monitor port. The IMC error point is not a good place to look for this signal because of the post-demodulation low pass filter (indeed, I didn't see any peak above the analyzer noise floor).

The nominal frequency was 11.066209 MHz, and I found that both peaks were simultaneously minimized by adjusting it to 11.066195 MHz, see Attachment #1. This corresponds to a length change of ~20 microns, which I think is totally reasonable. I guess the peaks can't be nulled completely because of imbalance in the positive and negative sidebands. 

Then, I checked for possible offsets at the IMC error point, by injecting a singal to the AO input of the IMC servo board (using the Siglent func gen), at ~300 Hz. I then looked at the peak height at the modulation frequency, and the second harmonic. The former should be minimized when the cavity is exactly on resonance, while the latter is proportional to the modulation depth at the audio frequency. I found that I had to tweak the MC offset voltage slider from the nominal value of 0V to 0.12 V to null the former peak, see Attachment #2. After accounting for the internal voltage division factor of 40, and using my calibration of the IMC error point as 13 kHz/V, this corresponds to a 40 Hz (~50 microns) offset from the true resonant point. Considering the cavity linewidth of ~4 kHz, I think this is a small detuning, and probably changes from lock to lock, or with time of day, temperature etc.

Conclusion: I think neither of these tests suggest that the IMC is to blame for the weirdness in the PRMI sensing, so the mystery continues.

Attachment 1: modFreq.pdf  206 kB  | Hide | Hide all
modFreq.pdf
Attachment 2: IMC_offset.pdf  157 kB  | Hide | Hide all
IMC_offset.pdf
ELOG V3.1.3-